OpenStudy (anonymous):

Question Already Answered. Just need someone to check. I will fan and medal anyone who does :D

OpenStudy (anonymous):

I will copy and paste the question

OpenStudy (anonymous):

what ish it?

OpenStudy (anonymous):

Solve. -1/8c = 2 A. –16 B. –4 C. 4 D. 16

OpenStudy (anonymous):

I got 16

OpenStudy (goose.):

Wouldn't it be -16? As it's-1/8c=2

OpenStudy (welshfella):

c = 2 / (-1/8) = 2 * -8

OpenStudy (anonymous):

-16 ^

OpenStudy (goose.):

You need to isolate your variable, so Multiply by -8 both sides.

OpenStudy (anonymous):

Ohh ok i know what i did wrong. I put 8 instead of -8

OpenStudy (anonymous):

Thanks guys :D

OpenStudy (goose.):

Haha, I assumed.

OpenStudy (goose.):

Gladly,

OpenStudy (welshfella):

-16 is correct